Problem 9-24 Critiquing a Report; Preparing a Performance Budget [LO9-1, LO9-2, LO9-3, LO9-4, LO9-6]

Exchange Corp. is a company that acts as a facilitator in tax-favored real estate swaps. Such swaps, know as 1031 exchanges, permit participants to avoid some or all of the capital gains taxes that would otherwise be due. The bookkeeper for the company has been asked to prepare a report for the company to help its owner/manager analyze performance. The first such report appears below:

Exchange Corp.
Analysis of Revenues and Costs
For the Month Ended May 31
Actual Unit Revenues
and Costs
Planning Budget
Unit Revenues
and Costs
Variances
Exchanges completed 30 25
Revenue $ 650 $ 750 $ 100 U
Expenses:
Legal and search fees 281 255 26 U
Office expenses 235 269 34 F
Equipment depreciation 25 30 5 F
Rent 75 90 15 F
Insurance 15 18 3 F
Total expense 631 662 31 F
Net operating income $ 19 $ 88 $ 69 U

Note that the revenues and costs in the above report are unit revenues and costs. For example, the average office expense is $269 per exchange completed on the planning budget; whereas, the average actual office expense is $235 per exchange completed.

Legal and search fees is a variable cost; office expenses is a mixed cost; and equipment depreciation, rent, and insurance are fixed costs. In the planning budget, the fixed component of office expenses was $5,250.

All of the company’s revenues come from fees collected when an exchange is completed.

Required:

1. Is the report prepared by the bookkeeper useful as a performance report?

2. Complete the performance report that would help the owner/manager assess the performance of the company in May.

Exchange Corp.
Flexible Budget Performance Report
For the Month Ended May 31
Actual Results Revenue and Spending Variances Flexible Budget Activity Variances Planning Budget
Exchanges completed 30 30 25
Revenue $19,500 $3,000 U $22,500 $3,750 F $18,750
Expenses:
Legal and search fees 8,430 780 U 7,650 1,275 U 6,375
Office expenses 7,050 ? U ? ? U 6,725
Equipment depreciation 750 0 None 750 0 None 750
Rent 2,250 0 None 2,250 0 None 2,250
Insurance 450 0 None 450 0 None 450
Total expense 18,930 ? U ? ? U 16,550
Net operating income $570 ? U ? ? F $2,200
Answer #1
1
NO, the report prepared by bookkeeper is not a useful performance report
2
Actual results Revenue and spending variances Flexible budget Activity variances Planning budget
Exchanges completed 30 30 25
Revenue 19500 3000 U 22500 3750 F 18750
Expenses:
Legal and search fees 8430 780 U 7650 1275 U 6375
Office expenses 7050 30 U 7020 295 U 6725
Equipment Depreciation 750 0 None 750 0 None 750
Rent 2250 0 None 2250 0 None 2250
Insurance 450 0 None 450 0 None 450
Total expense 18930 810 U 18120 1570 U 16550
Net operating income 570 3810 U 4380 2180 F 2200
Office expenses:
Variable component = (6725-5250)/25 = $59
Office expenses in flexible budget = 5250+(30*59) = $7020